OpenStudy (anonymous):

Solve 3x – 5y = 17 9x – 15y = –4

OpenStudy (anonymous):

solve first eq for x; x=(17+5y)/3. plug into x for second eq. 9((17+5y)/3)-15y=-4. then multiply out and simplify until you have aY=b, then Y=b/a. then plug this value of y into one of the original eqs. and solve for x

OpenStudy (anonymous):

3x – 5y = 17 9x – 15y = –4 Undefined solutions